a) Let d∈N , where d ≠1 and d is not divisible by the square of a prime.
Prove that N:Z[√d]→N∪{0}:N(a+b√d)=|a²-db²| satisfies the following properties for x,y∈Z[√d] :
i) N(x) = 0⇔x=0
ii) N(xy) = N(x)N(y)
iii) N(x) =1⇔ x is a unit
iv) N(x) is prime ⇒x is irreducible in Z[√d ] .
b) Prove or disprove that C≃ R as fields.
1
Expert's answer
2019-04-01T11:52:40-0400

a.

i) Let "N(a + \sqrt d b) = 0." It holds if and only if "|(a + \sqrt d b) (a - \sqrt d b)| = 0." It is true iff "(a + \sqrt d b) (a - \sqrt d b) = 0." Since "\mathbb{Z}[\sqrt d]" has no divisors of zero, we have either "a + \sqrt d b = 0" or "a - \sqrt d b = 0." And finally this holds if and only if "\sqrt d b = a = 0."

ii) "N(xy) = |xy \overline{xy}| = |x y \bar{x} \bar{y}| = |x \bar{x} y \bar{y}| = |x \bar{x}||y \bar{y}| = N(x)N(y)."

iii) If "N(x) = 1" and "x = a + \sqrt d b," we clearly have "\pm N(x) = (a + \sqrt d b) (a - \sqrt d b) = a^2 - db^2 = \pm 1." Then "x" is unit.

Now suppose that "x" is unit. Then we have "N(x)N(x^{-1}) = N(x x^{-1}) = N(1) = 1." But then the only possible value for "N(x)" is "1," since "N(x) \in \mathbb{N}."

iv) Let "N(x)" be prime and "x" be not irreducible. Then "x" can be represented as "x = ab," where "a" and "b" are not equal to one. Then "N(x) = N(ab) = N(a) N(b)." By iii) we have that "N(a)N(b) \not= 1." And "N(x) \not= 1." Contradiction.

b.

We will prove that "\mathbb{C}" and "\mathbb{R}" are not isomorphic.

Suppose the contrary: there exists an isomorphism "f: \mathbb{C} \to \mathbb{R}." Then we must have "f(1) = 1." Also "f(-1)^2 = f((-1)^2) = 1." Hence we have "f(i)^2 = f(i^2) = f(-1) = -1," contradiction.